# Entropic (Gromov) Wasserstein Flow Matching with GENOT

### v3

@@ A.1 Proofs of § 1
Proof.
This result follows directly from the construction of the loss. Suppose that $\mathcal{L}_{\mathrm{GENOT}}(\theta)=0$. By Fubini’s Theorem, this implies that:
$\begin{split}&\mathbb{E}_{X\sim\mu}\mathbb{E}_{t\sim\mathcal{U}([0,1]),Z\sim\rho,Y\sim\pi^{\star}_{\varepsilon}(\cdot|X)}[\|v_{t,\theta}\left([Z,Y]_{t}|X\right)-(Y-Z)\|_{2}^{2}]=0\,.\\ \textrm{i.e.}\quad&\mathbb{E}_{X\sim\mu}\mathbb{E}_{t\sim\mathcal{U}([0,1]),Z\sim\rho,Y\sim\pi^{\star}_{\varepsilon}(\cdot|X)}[\|v_{t,\theta}\left(tZ+(1-t)Y|X\right)-(Y-Z)\|_{2}^{2}]=0\,.\end{split}$ (6)
Since it is positive, the function
$\ell:\mathbf{x}\mapsto\mathbb{E}_{t\sim\mathcal{U}([0,1]),Z\sim\rho,Y\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})}[\|v_{t,\theta}\left(tZ+(1-t)Y|\mathbf{x}\right)-(Y-Z)\|_{2}^{2}]$
is therefore zero $\mu$-a.e., which means that for any $\mathbf{x}$ in the support of $\mu$, one has $\ell(\mathbf{x})=0$. Moreover, $\ell(\mathbf{x})$ is the CFM loss with source $\rho$ and target $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, applied to the vector field $v_{t,\theta}(\cdot|\mathbf{x})$. Therefore, $\ell(\mathbf{x})=0$ implies that $v_{t,\theta}(\cdot|\mathbf{x})$ generates a probability path between $\rho$ and $\pi^{\star}(\cdot|\mathbf{x})$ thanks to (Lipman et al., 2023, Theorem 1). This means that if $\phi_{t}(\cdot|\mathbf{x})$ is the flow induced by $v_{t,\theta}(\cdot|\mathbf{x})$, one has $\phi_{1}(\cdot|\mathbf{x})\sharp\rho=\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. Afterwards, if $X\sim\mu$, $Z\sim\rho$ and $Y=\phi_{1}(Z|X)$ denotes the solution of ODE (3), one has $Y\sim\pi^{\star}_{\varepsilon}(\cdot|X)$, so $(X,Y)\sim\pi^{\star}_{\varepsilon}$ by definition. ∎
See 3.2 We first remind the definition of a conditionally positive kernel, which is involved in point 2. of Prop. 3.2.
Definition A.1.
@@
We start with the linear case and consider $\pi^{\star}_{\varepsilon,\tau}$ solution of (LEOT). The result follows from duality. Indeed, from (Séjourné et al., 2023a, Prop. 2), one has the existence of the so-called entropic potentials $f^{\star}\in\mathcal{C}(\mathcal{X})$ and $g^{\star}\in C(\mathcal{Y})$ s.t.
@@
$\begin{split}&\eta(\mathbf{x})=\int_{\mathcal{Y}}\exp\left(\frac{f^{\star}(\mathbf{x})+g^{\star}(\mathbf{y})-c(\mathbf{x},\mathbf{y})}{\varepsilon}\right)\mathop{}\!\mathrm{d}\nu(\mathbf{y})\quad\textrm{and}\quad\xi(\mathbf{y})=\int_{\mathcal{X}}\exp\left(\frac{f^{\star}(\mathbf{x})+g^{\star}(\mathbf{y})-c(\mathbf{x},\mathbf{y})}{\varepsilon}\right)\mathop{}\!\mathrm{d}\mu(\mathbf{x})\,.\end{split}$ (8)
@@
$\begin{split}\min_{\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,+\varepsilon\mathrm{KL}^{\otimes}(\pi|\mu\otimes\nu)+\lambda_{1}\mathrm{KL}^{\otimes}(\pi_{1}|\mu)+\lambda_{2}\mathrm{KL}^{\otimes}(\pi_{2}|\nu),\end{split}$
@@
Step 2: Optimality in the balanced problem for the linear case. We now prove point 1, stating that if $\pi^{\star}_{\varepsilon,\tau}$ solves problem ULEOT between $\mu$ and $\nu$, then it solves problem (LEOT) between $\tilde{\mu}$ and $\tilde{\nu}$ for the same entropic regularization strength $\varepsilon$. It also follows from duality. We remind that that thanks to (Séjourné et al., 2023a, Prop. 2)
@@
with $f^{\star}\in\mathcal{C}(\mathcal{X})$ and $g^{\star}\in C(\mathcal{Y})$. Moreover, by (Nutz, , Theorem 4.2), such a decomposition is equivalent to the optimality in problem (LEOT). Therefore, $\pi^{\star}_{\varepsilon,\tau}$ is solves problem LEOT between its marginals $\tilde{\mu}$ and $\tilde{\nu}$, i.e.
@@
we get from (Séjourné et al., 2023a, Proposition 9), that $\pi^{\star}_{\varepsilon,\tau}$ solves:
@@
Since the costs are conditionally positive (or conditionally negative) kernels, (15) finally yields the desired result by applying (Séjourné et al., 2023b, Theorem 3):

### v4

@@ B.1 Existence of EOT couplings
We first remind the definition of a conditionally positive kernel, which is involved in point 2. of Prop. 3.1.
Definition B.1.
@@
We start with the linear case and consider $\pi^{\star}_{\varepsilon,\tau}$ solution of (LEOT). The result follows from duality. Indeed, from [80, Prop. 2], one has the existence of the so-called entropic potentials $f^{\star}\in\mathcal{C}(\mathcal{X})$ and $g^{\star}\in C(\mathcal{Y})$ s.t.
@@
$\begin{split}&\eta(\mathbf{x})=\int_{\mathcal{Y}}\exp\left(\frac{f^{\star}(\mathbf{x})+g^{\star}(\mathbf{y})-c(\mathbf{x},\mathbf{y})}{\varepsilon}\right)\mathop{}\!\mathrm{d}\nu(\mathbf{y})\\ \textrm{and}\quad&\xi(\mathbf{y})=\int_{\mathcal{X}}\exp\left(\frac{f^{\star}(\mathbf{x})+g^{\star}(\mathbf{y})-c(\mathbf{x},\mathbf{y})}{\varepsilon}\right)\mathop{}\!\mathrm{d}\mu(\mathbf{x})\,.\end{split}$ (11)
@@
$\begin{split}\min_{\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}&|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,\\ &+\varepsilon\mathrm{KL}^{\otimes}(\pi|\mu\otimes\nu)+\lambda_{1}\mathrm{KL}^{\otimes}(\pi_{1}|\mu)+\lambda_{2}\mathrm{KL}^{\otimes}(\pi_{2}|\nu),\end{split}$
@@
Step 2: Optimality in the balanced problem for the linear case. We now prove point 1, stating that if $\pi^{\star}_{\varepsilon,\tau}$ solves problem ULEOT between $\mu$ and $\nu$, then it solves problem (LEOT) between $\tilde{\mu}$ and $\tilde{\nu}$ for the same entropic regularization strength $\varepsilon$. It also follows from duality. We remind that that thanks to [80, Prop. 2]
@@
with $f^{\star}\in\mathcal{C}(\mathcal{X})$ and $g^{\star}\in C(\mathcal{Y})$. Moreover, by [59, Theorem 4.2], such a decomposition is equivalent to the optimality in problem (LEOT). Therefore, $\pi^{\star}_{\varepsilon,\tau}$ is solves problem LEOT between its marginals $\tilde{\mu}$ and $\tilde{\nu}$, i.e.
@@
we get from [80, Proposition 9], that $\pi^{\star}_{\varepsilon,\tau}$ solves:
@@
Since the costs are conditionally positive (or conditionally negative) kernels, (18) finally yields the desired result by applying [81, Theorem 3]: